OpenStudy (anonymous):

The demand function of a commodity is described by the exponential function P=10.5e^-0.1q where q is quantity demaded and TR=total revenue. Determine(i) The quantity for which the total revenue is maximized

OpenStudy (apoorvk):

so find the maxima of the function 'P' diferentiate P and equate to 0 ( i am assuming P is the TR) \[dP/dq = 10.5^{-o.1q} * (-0.1) = 0\] now this is zero at no point. so, more the 'q' that is demand, more the profit. thats what it should be i guess.

OpenStudy (anonymous):

Any one who can hwelp me i have an exam today

OpenStudy (anonymous):

\[TR = Pq\]\[TR = \left( 10.5 \right)qe^{-0.1q}\]maximize the total revenue by using the first derivative\[\frac{dTR}{dq}=0\]

OpenStudy (anonymous):

How do i diffrentiate an Exponential Exraven

OpenStudy (anonymous):

I will show you how to differentiate TR\[TR = \left( 10.5 \right)qe^{-0.1q}\]use the product rule\[u = 10.5q\]\[u' = 10.5\]\[v = e^{-0.1q}\]\[v' = -0.1e^{-0.1q}\]\[TR' = u'v + uv'\]\[TR' = 10.5e^{-0.1q} - 1.05qe^{-0.1q}\]

OpenStudy (anonymous):

to differentiate an exponential, recall that\[\frac{d}{dx}\left( e^{x} \right)=e^{x}\]

OpenStudy (anonymous):

Thaxs a million tymes..

OpenStudy (anonymous):

Extraven when i diffrentiate i get Q=10 am i right?

OpenStudy (anonymous):

firm producing hokey sticks has a production function given by Q=2√(K.L) Where: K= Capital inputs L= Labour inputs In the short run, the firm amounts of capital equipments is Fixed at K=100. The rental rate for capital is V=1.00 Shillings and the wage rate for L=4.00 shillings. Calculate the firms short run costs functions as well as the short run average costs Functions. What is the firms short run marginal costs.